### v1

# Beware of the Classical Benchmark Instances for the Traveling Salesman Problem with Time Windows

Abstract
We propose a simple and exact informed search method for the Traveling Salesman Problem with Time Windows and Makespan objective (TSPTW-M) that solves all instances of the classical benchmark with $50$ or more customers in less than ten seconds each. Applying this algorithm as an off-the-shelf method, we also solve all but one of these instances for the Duration objective. Our main conclusion is that these instances should no longer be employed for evaluating the TSPTW-M and its Duration variant: they can be “hacked” to yield results that seem outstanding at first sight. Additionally, caution is advised when designing hard training sets for machine learning algorithms.
Keywords: traveling salesman problem with time windows, makespan objective, duration objective, informed search, benchmark instances, training datasets.
1 Introduction
In a Traveling Salesman Problem with Time Windows (TSPTW), a vehicle must visit a set of customers within their predefined time windows. The vehicle cannot arrive late to a customer, although it can arrive early and wait until the beginning of its time window. The journey begins at a depot to which the vehicle must return after visiting all the customers. Several variants of the TSPTW are defined, depending on the objective function. Three of the most studied variants involve minimizing the time to complete the route (TSPTW-M), the duration of the route (TSPTW-D), or the travel time of the vehicle, disregarding waiting times (TSPTW-TT). From here on, we use the acronym TSPTW to refer to one of these three variants.
Research on the TSPTW began over forty years ago with foundational papers, such as ChristofidesMingozziToth1981 and Baker1983 for the TSPTW-M, Savelsbergh1985 for the TSPTW-TT, and Savelsbergh1992 for the TSPTW-D. Since then, numerous algorithms employing various solving strategies have been proposed for these three problem variants.
For exact approaches, significant efforts have focused on developing strong formulations suited for branch-and-bound and branch-and-cut methods (e.g. Baker1983; LangevinDesrochersDesrosiersGelinasSoumis1993; AscheuerFischettiGroetschel2001; KaraKocAltiparmakDengiz2013), including time-expanded networks (e.g. DashGuenluekLodiTramontani2012; BolandHewittVuSavelsbergh2017), dynamic programming (both pure and incomplete) (e.g. DumasDesrosiersGelinasSolomon1995; MingozziBiancoRicciardelli1997; BalasSimonetti2001), and anytime informed search methods (e.g. FontaineDibangoyeSolnon2023; KuroiwaBeck2023). Other methods include dynamic programming within column generation and Lagrangian relaxation (e.g. BaldacciMingozziRoberti2012; TilkIrnich2017; Lera-RomeroMirandaBrontSoulignac2022), constraint programming (e.g. PesantGendreauPotvinRousseau1998; FocacciLodiMilano2002), and more recently, (multivalued) decision diagrams (e.g. GillardCoppeSchausCire2021; RudichCappartRousseau2023; CoppeGillardSchaus2024).
In terms of heuristic and metaheuristic approaches, in addition to the pioneering works by Savelsbergh1985 and Savelsbergh1992, notable contributions include heuristic methods by GendreauHertzLaporteStan1998; Calvo2000; Helsgaun2017, tabu search by CarltonBarnes1996, compressed annealing by OhlmannThomas2007, and ant colony optimization (ACO) along with its Beam-ACO extensions by FavarettoMorettiPellegrini2006; Lopez-IbanezBlum2010; Lopez-IbanezBlumOhlmannThomas2013. Other significant approaches include general variable neighborhood search (e.g. daSilvaUrrutia2010; MladenovicTodosijevicUrosevic2013; AmgharCordeauGendron2019; YeBartoliniSchneider2024), iterated maximum large neighborhood search by Pralet2023, and more recently, machine learning-based solvers (e.g. CappartMoisanRousseauPremont-SchwarzCire2021; KoolHoofGromichoWelling2022; ZhengHeZhouJinLi2023; BiMaZhouSongCaoWuZhang2024; ChenGongChenLiuWangYuZhang2025; li2025lmasklearnsolveconstrained), among others.
Despite the many techniques developed for the TSPTW, the performance of most algorithms is evaluated on a subset of the classical benchmark instances compiled by López-Ibáñez and Blum in https://lopez-ibanez.eu/tsptw-instances, regardless of the objective function under consideration. The common reasoning is that the validity of an instance depends more on the space of feasible routes than on the objective function. Many machine learning algorithms replicate the generation process of some of these instances to create the large datasets of “hard” instances needed for training purposes (see e.g. CappartMoisanRousseauPremont-SchwarzCire2021; KoolHoofGromichoWelling2022; BiMaZhouSongCaoWuZhang2024; li2025lmasklearnsolveconstrained). A recent exception to this rule for evaluating the TSPTW is Fontaine2023, who designs hard instances with few customers based on a benchmark by RifkiChiabautSolnon2020.
From the perspective of exact solvers, the classical benchmark is nearing the end of its useful life. BaldacciMingozziRoberti2012 solved all but one of the instances for the TSPTW-TT in at most one hour; RudichCappartRousseau2023; FontaineDibangoyeSolnon2023 jointly solved all but five instances for the TSPTW-M in at most one hour; and Lera-RomeroMirandaBrontSoulignac2022 solved all but fourteen of the considered instances for the TSPTW-D in at most three hours. A natural question then arises: how will we develop the next generation of benchmarking instances? daSilvaUrrutia2010 addressed this question by creating a benchmark of large instances with $400$ customers, using similar methods to those employed for the classical instances. Interestingly, FontaineDibangoyeSolnon2023 solved the benchmark by daSilvaUrrutia2010 in a few minutes, even though it is unable to solve some instances by Fontaine2023; RifkiChiabautSolnon2020 with $30$ customers within an hour.
The idea that the hardest instances for the TSPTW are those with loose time windows is widely accepted in the literature (see e.g. DumasDesrosiersGelinasSolomon1995). What sets the benchmark proposed by Fontaine2023 apart is its use of a parameter, $\beta$, to systematically widen the time windows, as suggested by AriglianoGhianiGriecoGuerrieroPlana2019 for time-dependent problems. In a recent study, RifkiSolnon2025 investigate the effects of $\beta$ and another parameter, $\alpha$, on the feasibility and hardness of Euclidean instances for the TSPTW-M. Their main finding regarding feasibility is the presence of a pronounced phase transition between the infeasible and feasible regions. On the topic of hardness, they empirically demonstrate that the exact method proposed by FontaineDibangoyeSolnon2023 benefits from tight time windows. As the authors note, their analysis is crucial for the development of next-generation of benchmark instances and for training machine learning algorithms that requires datasets with varying levels of difficulty. More recently, ChenGongChenLiuWangYuZhang2025 have questioned the time window generation processes used in previous machine learning studies, suggesting that they result in weak instances that are too easy to solve.
1.1 Our contributions
The main purpose of this article is to reinforce the conclusions of RifkiSolnon2025 and ChenGongChenLiuWangYuZhang2025, demonstrating that the classical benchmark instances with $50$ or more customers, as well as the benchmark instances by daSilvaUrrutia2010, should no longer be used to evaluate the TSPTW-M. This is because these instances can be solved by a simple exact algorithm in less than ten seconds each. Indeed, any method that includes a variant of our solver as a preprocessing step could yield misleading results that appear outstanding at first glance, regardless of how well does the algorithm performs after this preprocessing phase. Our method works by repeatedly running a simple informed search in a backward direction, prioritizing partial routes that reach the end depot earlier. Despite its simplicity, it is surprising that the algorithm can solve all the large instances, including the four that remained open in previous studies. However, the algorithm has serious limitations: it fails to solve some classical benchmark instances with fewer than $50$ customers and cannot solve any instance in the benchmark by Fontaine2023; RifkiChiabautSolnon2020 with 30 or 40 customers and loose time windows.
Regarding the TSPTW-D, we show that using our solver for the TSPTW-M as an off-the-shelf method is sufficient to solve all but one of the classical benchmark instances with $50$ or more customers, including the fourteen instances that remained unsolved in the study by Lera-RomeroMirandaBrontSoulignac2022. This time, all instances are solved in less than $30$ minutes each. These results further reinforce the notion that the classical benchmark instances are particularly easy for our informed search method and, therefore, they should not be used to evaluate the efficiency of algorithms for the TSPTW-D either.
Although our main conclusion is that the classical benchmark instances should be avoided when evaluating the TSPTW-M and the TSPTW-D, we believe that caution is also required when evaluating the TSPTW-TT. Moreover, we argue that our results should be considered when designing new benchmark sets or training datasets for machine learning methods. Neglecting to do so could lead to biased conclusions about the efficiency and effectiveness of the solvers (ChenGongChenLiuWangYuZhang2025). The ideas proposed by AriglianoGhianiGriecoGuerrieroPlana2019, Fontaine2023, and RifkiSolnon2025, which use a time window tightness parameter $\beta$ to widen the time windows, provide a solid foundation for generating harder benchmark instances–even with fewer customers.
The fact that changing $\beta$ yields harder instances certainly helps researchers evaluate the efficiency of their methods, but it does not imply that the synthetic time windows generated are significant for real-life situations. For example, Ascheuer1996 discuss a real-life problem where tight time windows are designed to ensure that jobs are completed promptly in an online setting. The short question, whose answer remains unclear to us, is whether the classical benchmark instances are truly representative of real-world problems. An affirmative answer could suggest that our algorithm, while too simplistic as a general purpose approach, might still be valid as a preprocessing step for filtering out easy instances before applying more sophisticated techniques.
2 Problem Statement
Throughout this article, we write $[j]=[0,j]$, $i,j]\!]=[i,j]\cap\mathbb{N}$, and [[j]]=[[0,j]][\![j]\!]=[\![0,j]\!] for $i,j\in\mathbb{R}$. Consider a transport network described by a complete digraph $D$ with vertex set [[n+1]][\![n+1]\!]. Vertices 0 and $n+1$ represent the start and end depots, respectively, whereas the vertices in [[1,n]][\![1,n]\!] correspond to customers. Each arc $v\to w$ of $D$ has a travel time $\tau(v,w)$ representing the time required to travel from $v$ to $w$. Additionally, each vertex $v\in[\![n+1]\!]$ has a nonempty time window $[a(v),b(v)]$ in which $v$ must be visited. We assume that $a(v),b(v)]\subseteq[a(0),b(0)]=[a(n+1),b(n+1)]=[T]$ for $v\in[\![1,n]\!]$, where $T$ is the planning horizon.
A partial route is a nonempty sequence of vertices in $D$. A partial route $R=\langle v_{0},\ldots v_{k}\rangle$ is a route if $v_{0}=0$ and $v_{k}=k=n+1$, while it is elementary when $v_{i}\neq v_{j}$ for $\in[\![k]\!], iji\neq j$. Thus, $R$ is an elementary route if and only if it visits each customer in $D$ exactly once.
For $i\in[\![k]\!]$ and $t\in[T]$, we can compute the earliest arrival time to $v_{i}$ when the vehicle traverses $R$ departing from $v_{0}$ at time $t$ using the following recurrence that takes waiting times into account:
$\delta(R,t,i)=\begin{cases}\max\{t,a(v_{0})\}&\text{if }i=0,\\ \max\{\delta(R,t,i-1)+\tau(v_{i-1},v_{i}),a(v_{i})\}&\text{otherwise.}\end{cases}$
The partial route $R$ is feasible for time $t$ if $t\geq a(v_{0})$ and $\delta(R,t,i)\leq b(v_{i})$ for every $i\in[\![k]\!]$. Let $\delta(R,t)=\delta(R,t,k)$ if $R$ is feasible, and $\delta(R,t)=\infty$ otherwise. We refer to $\delta(R,t)$ as the (earliest feasible) arrival time of $R$ departing at time $t$. The makespan of $R$ is $\delta(R)=\delta(R,a(v_{0}))$, while its duration is $\Delta(R)=\min\{\delta(R,t)-t\mid t\in[T]\}$. The goal of the TSPTW-M (resp. TSPTW-D) is to find an elementary route with minimum makespan (resp. duration). Observe that the makespan (resp. duration) of such optimum routes is infinity when no elementary route is feasible for time 0.
Sometimes it is convenient to know the departure time of the vehicle to arrive the last node of a partial route $R$ at a given time $t\in[T]$. We let $\delta^{-1}(R,t)$ be the maximum $t^{\prime}\in[T]$ such that $\delta(R,t^{\prime})=t$ ($t^{\prime}=$-$\infty$ if no such $t^{\prime}$ exists). We refer to $\delta^{-1}(R,t)$ as the (latest feasible) departure time of $R$ arriving at time $t$.
2.1 Benchmark instances
In this article, we consider $10$ sets of benchmark instances, comprising a total of $1337$ instances. The first $7$ sets contain $467$ instances and correspond to the classical instances compiled by López-Ibáñez and Blum, available at https://lopez-ibanez.eu/tsptw-instances. As explained in Section 1, the efficiency of most algorithms solving TSPTWs is evaluated on a subset of these instances. Following Fontaine2023, we refer to the individual benchmarks as Asc (Ascheuer1996), Dum (DumasDesrosiersGelinasSolomon1995), Gen (GendreauHertzLaporteStan1998), Lan (LangevinDesrochersDesrosiersGelinasSoumis1993), Ohl (OhlmannThomas2007), Pes (PesantGendreauPotvinRousseau1998), and Pot (PotvinBengio1996). The travel times in Ohl are obtained by rounding the original distances to the nearest integer. The eighth benchmark, called Ohl-D, contains the same $25$ instances as Ohl but with travel times rounded to the to the nearest tenth. This benchmark was considered in the context of the TSPTW-D. The ninth set, called DaS, comprises the $125$ instances from daSilvaUrrutia2010, while the last set, called Rif, contains the remaining $720$ instances from Fontaine2023, based on a benchmark by RifkiChiabautSolnon2020.
To provide some context for the results discussed in the subsequent sections, we briefly review the origins of these benchmark sets. Our primary focus is on the time windows, as their distributions appear to be more significant than the number of customers (see e.g. RifkiSolnon2025). We also note that many training sets for machine learning methods are generated using procedures similar to those employed in the creation of Lan, Dum, Gen, Ohl, and DaS (e.g. CappartMoisanRousseauPremont-SchwarzCire2021; KoolHoofGromichoWelling2022; BiMaZhouSongCaoWuZhang2024; li2025lmasklearnsolveconstrained).
  • Lan instances distribute the locations of a depot and $n\in\{20,40,60\}$ customers uniformly in $\sigma]\times[\sigma]$ with $\sigma=100$. The Euclidean distance, rounded to the nearest tenth, is used as the travel time. Time windows are set by computing a second nearest-neighbor route $R=\langle v_{1},\ldots,v_{n+1}\rangle$, and assigning $a(v_{i})=\delta(R,0,i)-\mathcal{U}[0,\omega/2]$ and $b(v_{i})=\delta(R,0,i)+\mathcal{U}[0,\omega/2]$ for a width $\omega\in\{40,60,80\}$. These instances are now considered trivial for the TSPTW-M.
  • Dum instances were created using the same procedure as Lan, but with $\sigma=50$ for $n\in\{20,40,60\}$ and $\omega\in\{20,40,60,80,100\}$; $n=80$ and $\omega\in\{20,40,60,80\}$; $n\in\{100,150\}$ and $\omega\in\{20,40,60\}$; and $n=200$ and $\omega\in\{20,40\}$. Quoting DumasDesrosiersGelinasSolomon1995, “for narrow widths, [our algorithm] is less than exponential […,] a 250-node problem with $\ell=20$ is easy to solve in less than 10 seconds.”
  • Gen extends the benchmark by Dum for $n\in\{20,40,60,80,100\}$ and $\ell\in\{60,80,\ldots,200\}$. In their words, “[Our algorithm] would appear to be slower […] on instances with narrow time windows. As noted by Dumas et al., the running time of their exact algorithm increases exponentially with time window width. [We generated our instances] to test the performance […] on instances with wide time windows.”
  • Ohl also extends Dum for $n=150$ and $\ell\in\{120,140,160\}$, and for $n=200$ and $\ell\in\{120,140\}$. Their justification is that their instances “show a solution method’s ability to not only cope with wide time windows, but also with large numbers of customers.”
  • DaS was designed to “overcome the limitations of the other test sets” as daSilvaUrrutia2010 needed “instances with more than 200 customers and wider time windows to benchmark and compare algorithms.” Additionally, they believed that the method by DumasDesrosiersGelinasSolomon1995 was “a biased way to generate instances since it does not reflect the real cases. Moreover, there is no challenge to build a feasible solution because the second-nearest neighbor tour already is a feasible solution.” To address these concerns, they generate random routes to assign the time windows, with $n\in\{200,250,300,350,400\}$, $\sigma=100$, and $\omega\in\{100,200,300,400,500\}$.
  • Pot and Pes instances were obtained by extracting the route of a vehicle from a solution to Solomon’s RC2 instances for the vehicle routing problem with time windows. They are considered hard but have at most 45 customers.
  • Asc instances are derived from a real-world application that minimizes the unloaded travel time of a stacker crane in an online setting. The number of customers ranges from $10$ to $231$, while the travel time windows are tight to ensure jobs are not delayed for too long.
  • RifkiChiabautSolnon2020 designed a benchmark set for time-dependent routing problems with travel times computed from shortest paths in the road network of Lyon, using realistic traffic simulations built from real-world data. To create Rif, Fontaine2023 removed the time dependency from the instances with $n\in\{20,30,40\}$. Regarding time windows, Fontaine2023 follows the model presented by AriglianoGhianiGriecoGuerrieroPlana2019 for time-dependent instances, setting $a(v_{i})=(\beta\delta(R,0,i)-40)$ and $b(v_{i})=\delta(R,0,i)+40$ for a random route $R=\langle v_{1},\ldots,v_{n+1}\rangle$ and $\beta\in\{0,0.25,0.5,1\}$. Here, $\beta\in[1]$ represents the tightness of the time windows, from the loosest 0 to the tightest $1$.
3 A Simple Solver for the TSPTW-M
The solver for the TSPTW-M builds upon the best-first search algorithm depicted in Algorithm 1. Given $t_{0},ub\in[T]$, Algorithm 1 determines whether an elementary route $R$ in the transport network $D$ is feasible and arrives earlier than $ub$ when departing at time $t_{0}$, i.e., whether $\delta(R,t_{0})\leq ub$. If $t_{0}=0$, then the makespan of $R$ is less than $ub$.
For convenience, for every partial route $R$ starting at vertex $v$, we define: $\delta_{\rm m}(R)=\delta(R,\max\{t_{0},a(v)\})$ as the earliest arrival time when the vehicle leaves the depot at time $t_{0}$, and $\delta^{-1}_{\rm m}(R)=\delta^{-1}(R,ub)$ as the latest departure time when the vehicle cannot arrive at the end depot after time $ub$.
In essence, Algorithm 1 is an informed search procedure that traverses a tree of partial routes $\mathcal{T}$ in a backward direction. The root of $\mathcal{T}$ is the partial route $\langle n+1\rangle$, which only visits the end depot, while the children of a partial route $R$ are all the elementary partial routes $\langle w\rangle+R$ obtained by prepending a vertex $w\not\in V(R)$ to $R$. Here, $V(R)$ denotes the set of vertices set already visited by $R$.
Within Algorithm 1, $\mathcal{R}$ maintains the unprocessed nodes of $\mathcal{T}$, which have already had their parent processed (Steps 3 and 14). Since $\mathcal{T}$ contains $\Theta(n!)$ nodes, two techniques are employed in Step 14 to prevent the exploration of the entire tree. First, a set of unreachable vertices $U(w,\delta^{-1}_{\rm m}(R))$ is computed for every partial route $R$ that starts at vertex $w$, as described in Section 3.1. By construction, if $U(w,\delta^{-1}_{\rm m}(R))\not\subseteq V(R)$, then no elementary route $R^{\prime}$ with suffix $R$ and $\delta(R^{\prime})<ub$ exists, and therefore the subtree of $\mathcal{T}$ rooted at $R$ can be discarded. Similarly, if $\delta_{\rm m}(R)\geq ub$, then $\delta_{\rm m}(R^{\prime})\geq ub$ for every route $R^{\prime}$ with suffix $R$, and so this subtree can also be pruned. These pruning strategies ensure that any partial route $R$ obtained in Step 8 is feasible and has $\delta_{\rm m}(R)<ub$.
Following FontaineDibangoyeSolnon2023, each iteration of the main loop (Steps 414) processes one route $R$ of length $k$, if one exists, for every $k\in[\![n+1]\!]$ (Steps 514). All extensions of $R$ in Step 14 will have length $k+1$, and thus can be processed within the same iteration of the main loop. The goal is to compute the output route as early as possible. To guide the search, a heuristic is applied to decide which partial route $R$ of length $k$ to process next (Step 7).
Note that $ub-\delta_{\rm m}(R)$ represents an upper bound on the time that can still be saved by any route with suffix $R$. In the extreme case where $\delta_{\rm m}(R)\geq ub$, no improvement is possible, and $R$ is discarded (Step 14). Maximizing $ub-\delta_{\rm m}(R)$ maximizes the opportunity for the vehicle to wait at previous customers, if necessary. In event of ties, Step 7 minimizes $\delta^{-1}(R,\delta_{\rm m}(R))$ to maximize the earliest departure time of $R$ that yields no waiting due to time windows. The goal is to allow the vehicle to wait at the previous (yet unknown) customers without significantly affecting the earliest arrival time.
Finally, Algorithm 1 avoids extending a partial route $R$ if it can prove that it is $ub$-dominated. Formally, $R$ is $ub$-dominated by a partial route $Q$ of length $|R|$ if, for every route $R^{\prime}$ with suffix $R$, either the earliest arrival time of $R^{\prime}$ is $\delta_{\rm m}(R^{\prime})\geq ub$, or there exists a route $Q^{\prime}$ with suffix $Q$ whose earliest arrival time is $\delta_{\rm m}(Q^{\prime})<ub$. The extension of $R$ can be avoided when it is $ub$-dominated by an extended partial route $Q\neq R$. Algorithm 1 applies the following rule to discard $ub$-dominated routes.
Proposition 1.
Let $R$ be a partial route from vertex $v$ to the end depot $n+1$ with $\delta^{-1}_{\rm m}(R)\geq t_{0}$. If there exists a partial route $Q\neq R$ from $v$ to $n+1$ such that its earliest arrival time is $\delta_{\rm m}(Q)<ub$, it visits the vertices $V(Q)=V(R)$, its latest departure is $\delta^{-1}_{\rm m}(Q)\geq\delta^{-1}_{\rm m}(R)$ and one of the following holds:
  1. 1.
    its latest departure is $\delta^{-1}_{\rm m}(Q)>\delta^{-1}_{\rm m}(R)$, or
  2. 2.
    $\delta(R,\delta^{-1}_{\rm m}(R))=ub$ (i.e., $R$ cannot arrive earlier than $ub$ when departing as late as possible), or
  3. 3.
    $\delta(Q,\delta^{-1}_{\rm m}(Q))<ub$ (i.e., $Q$ arrives earlier than $ub$ when departing as late as possible).
then $R$ is $ub$-dominated by $Q$.
Steps 913 apply Proposition 1 to determine if $R$ is $ub$-dominated. A table is maintained, indexed by a vertex set $V$ and a vertex $v$, to store the greatest $\delta^{-1}_{\rm m}(Q)$ among every processed route $Q$ with $V(Q)=V$ and first vertex $v$. The table also stores $\delta(Q,\delta^{-1}_{\rm m}(Q))$; note that $\delta_{\rm m}(Q)<ub$ follows by invariant. Because of the search heuristic, Algorithm 1 can process an $ub$-dominated partial route before extending any of its dominators. We note that Algorithm 1 could discard all the routes minimizing the makespan because some of their prefixes are $ub$-dominated. This explains why Algorithm 1 is only a decision method rather than an optimization method, which justifies the decision to stop as soon as a route with $\delta_{\rm m}(R)<ub$ is found.
Algorithm 1 Backward Best First Search Labeling
1:a transport network $D$, a departure time $t_{0}\in[T]$ and a latest arrival time $ub\in[T]$.
2:a route $R$ with $\delta_{\rm m}(R)<ub$ or a message that no such $R$ exists.
3:let $\mathcal{R}=\{\langle n+1\rangle\}$ be a family of partial routes
4:while $R$ is nonempty:
5:  for all $k\in[\![n+1]\!]$:
6:   if $\mathcal{R}$ has no route of length $k$: continue
7:   Remove from $\mathcal{R}$ a partial route $R$ of length $k$ minimizing $\langle\delta_{\rm m}(R),\delta^{-1}(R,\delta_{\rm m}(R))\rangle$
8:   if $k=n+1$: output $R$ and halt
9:   let $v$ be the first vertex in $R$ and $(d,a):=\operatorname{best}(V(R),v)$
10:   if $\delta^{-1}_{\rm m}(R)<d$: continue
11:   if $\delta^{-1}_{\rm m}(R)=d$ and $a<ub$: continue
12:   if $\delta^{-1}_{\rm m}(R)=d$ and $\delta(R,\delta^{-1}_{\rm m}(R))=ub$: continue;
13:   else: let $\operatorname{best}(V(R),v)=\langle\delta^{-1}_{\rm m}(R),\delta(R,\delta^{-1}_{\rm m}(R))\rangle$
14:   for all $w\in[\![n]\!]-V(R)$: insert $R^{\prime}=\langle w\rangle+R$ into $\mathcal{R}$ if $\delta_{\rm m}(R^{\prime})<ub$ and $U(w,\delta^{-1}_{\rm m}(R^{\prime}))\subseteq V(R^{\prime})$
15:output “no route $R$ exists with $\delta_{\rm m}(R)<ub$
To solve the TSPTW-M, we repeatedly execute Algorithm 1 as described in Algorithm 2. The output of this process is the route that minimizes the earliest arrival time, $\delta(R,t_{0})$, for a given departure time $t_{0}\in[T]$.
Algorithm 2 Basic solver for the TSPTW-M
1:a transport network $D$ and a departure time $t_{0}\in[T]$
2:a route $R$ minimizing $\delta_{\rm m}(R)<T$ or a message no such route exists.
3:let $R^{*}=\emptyset$ and $T^{\prime}=T+1$.
4:Compute the unreachable function for Step 14 as in Section 3.1
5:while Algorithm 1 with input $D$, $t_{0}$, and $ub=T^{\prime}$ outputs a route $R$:
6:  let $T^{\prime}=\delta_{\rm m}(R)$ and $R^{*}=R$
7:if $R^{*}\neq\emptyset$: output $R^{*}$, else output “the desired route does not exist”.
Implementation details.
Although our implementation of Algorithm 1 is not particularly efficient, we have taken measures to improve its running time. Each partial route $R$ is implemented with a label $\ell(R)$ that uses $O(1)$ bits. The label $\ell(R)$ is a tuple containing the following information:
  • The first vertex $v$ of $R$,
  • A bitset representing the set of visited vertices $V(R)$,
  • The latest departure time $\delta^{-1}_{\rm m}(R)$,
  • The earliest arrival time $\delta_{\rm m}(R)$,
  • The other limiting times $\delta(R,\delta^{-1}_{\rm m}(R,ub))$ and $\delta^{-1}(R,\delta_{\rm m}(R))$, and
  • A pointer to the label representing its parent in $\mathcal{T}$.
By traversing the branch of $\mathcal{T}$ from node $R$ using the parents pointers, we can obtain all the vertices in $R$ in $O(n)$ time. Each time $R$ is extended into a a partial route $R^{\prime}$ in Step 14, the times $\delta(R^{\prime},\delta^{-1}_{\rm m}(R^{\prime}))$ and $\delta^{-1}(R^{\prime},\delta_{\rm m}(R^{\prime}))$ are computed in $O(n)$ time. Additionally, $U(w,\delta^{-1}_{\rm m}(R^{\prime}))$ is computed in $O(n)$ time, as discussed in Section 3.1.
The family of routes $\mathcal{R}$ is implemented using a priority queue that holds all the unprocessed routes of length $k$, for every $k\in[\![n+1]\!]$. Finally, we maintain a hash table $\operatorname{best}[v]$ for each $v\in[\![n+1]\!]$, indexed by $V(R)$, which is used in Steps 9 and 13.
3.1 The Unreachable Function
As illustrated in Algorithm 2 (Step 4), we apply a preprocessing step to compute an unreachable function $U\colon[\![n+1]\!]\times[T]\to 2^{[\![n+1]\!]}$, which is later used by Algorithm 1 to prune some branches of the tree $\mathcal{T}$ of partial routes. For each vertex $v\in[\![n+1]\!]$ and time $t\in[T]$, the function $U(v,t)$ returns a set of vertices that cannot appear before $v$ in any elementary and feasible route $R$ that visits $v$ at a time $t^{\prime}\leq t$. By construction, $U(v,t)\subseteq U(v,t^{\prime})$, so $U$ is implemented by storing $U(v,t)$ only for those times $t\in[T]$ where the function changes. As a result, each $U(v,\bullet)$ has $O(n)$ values, which allows for queries to be answered in $O(n)$ time.
The computation of $U$ follows these steps. First, we run an all-pairs shortest path algorithm twice to obtain the following functions:
[…]
for every pair of vertices $v,w$ in $D$. Here, $\operatorname{EAT}(w,v)$ represents the earliest arrival time at $v$ after visiting $w$, and $\operatorname{LDT}(w,v)$ represents the latest departure time from $w$ before visiting $v$.
Next, we compute the precedence relation $\prec$, where $v\prec w$ if and only if $\operatorname{EAT}(w,v)=\infty$ or there exists a third vertex $z\in[\![n+1]\!]-\{v,w\}$ such that:
$\operatorname{EAT}(z,w)>\operatorname{LDT}(w,v),\quad\operatorname{EAT}(w,v)>\operatorname{LDT}(v,z),\quad\text{and}\quad\operatorname{EAT}(w,z)>\operatorname{LDT}(z,v).$
This relation ensures that no feasible route can visit $w$ before $v$ when $v\prec w$. Therefore, for every $t\in\mathbb{R}$, we set $w\in U(v,t)$ if $v\prec w$. Otherwise, we set $w\in U(v,t)$ if and only if $t<\operatorname{EAT}(w,v)$.
Once the precedence relation $\prec$ is computed, we update the time windows by setting:
$a(w):=\max\{a(w),\operatorname{EAT}(v,w)\}\quad\text{and}\quad b(v):=\min\{b(v),\operatorname{LDT}(v,w)\}$
for every pair of vertices $v,w\in[\![n+1]\!]$ such that $v\prec w$. If any time window is modified, we recompute $U$ and repeat the process until no further changes are made.
3.2 Solving the integer TSPTW-D
Recall that the TSPTW-D consist of finding an elementary route $R$ with minimum duration $\Delta(R)=\min\{\delta(R,t)-t\mid t\in[T]\}$ for a given transport network $D$. Clearly, if the travel times and all the time windows in $D$ are integers, then a solution to the TSPTW-D can be obtained by running a solver for the TSPTW-M $O(T)$ times. This is the approach Algorithm 3 uses to solve the TSPTW-D.
Specifically, Algorithm 3 maintains a sliding window [[p,q]][\![p,q]\!], a route $R$ that minimizes $\delta(R,p-1)=q$, and a route $R^{*}$ with a time $t^{*}$ that minimizes $\delta(R^{*},t^{*})-t^{*}$ for $t^{*}\in[\![p-1]\!]$. Steps 35 initialize these variables. Step 5 also computes $\operatorname{last}_{p}$, which is the latest of the departure times of all the elementary routes in $D$. Thus, when $p>\operatorname{last}_{p}$, $R^{*}$ is a route with minimum duration.
Throughout the main loop (Steps 611) the route $R$ is updated into a new route that minimizes $\delta(R,p)\geq q$. Note that $\delta(R,p)=q$ when Step 9 is skipped. Then, $q$, $R^{*}$, and $t^{*}$ are updated accordingly in Steps 9, 11, and 12.
Algorithm 3 Sliding window solver for the integer TSPTW-D
1:a transport network $D$ in which $\tau$ and the time windows are integer
2:a route $R^{*}$ and a time $t^{*}$ minimizing $\Delta(R^{*})<T$ or a message that $D$ is infeasible.
3:let $R$ be a route minimizing $\delta(R)$; halt with outputDD is infeasible” if $\delta(R)=\infty$
4:Heuristically find a route $R^{*}$ and a time $t^{*}$ with $\delta(R^{*},t^{*})=\delta(R)$ maximizing $t^{*}$
5:let $p=t^{*}+1$, $q=\delta(R^{*})$, and $\operatorname{last}_{p}=\max\{\delta^{-1}(R,T)\mid R\text{ is an elementary route}\}$
6:while $p\leq\operatorname{last}_{p}$:
7:  Heuristically update $R$ to minimize $\delta(R,p)$
8:  if $\delta(R,p)>q$:
9:   let $R$ be a route minimizing $\delta(R,p)$ and let $q=\delta(R,p)$
10:  Heuristically update $R$ and $p$ to maximize $p$ with $\delta(R,p)=q$
11:  if $\delta(R^{*},t^{*})-t^{*}>\delta(R,p)-p$: let $R^{*}=R$ and $t^{*}=p$
12:  let $p=p+1$
13:output $R^{*}$ and $t^{*}$
Regarding the implementation of Algorithm 3, we run Algorithm 2 with $t_{0}=0$ and $ub=T+1$ for Step 3. Similarly, we invoke Algorithm 2 with $t_{0}=p$ and $ub=\min\{\operatorname{last}_{q}+\Delta(R^{*}),T+1\}$ for Step 9. Let $D^{-1}$ be the reverse transport network with vertex set [[n+1]][\![n+1]\!], where the travel time of an arc $v\to w$ is $\tau(w,v)$, and the time window of a vertex $v$ is [Tb(v),Ta(v)][T-b(v),T-a(v)]. In $D^{-1}$, the start depot is $n+1$ and the end depot is 0. To initialize $\operatorname{last}_{p}$ in Step 5, we run Algorithm 2 on $D^{-1}$ with $t_{0}=0$ and $ub=T-t^{*}$. If a route $R^{-1}$ in $D^{-1}$ is found, then $\operatorname{last}_{p}=T-\delta(R^{-1},0)$; otherwise, $\operatorname{last}_{p}=t^{*}$. Finally, we use a local search heuristic for Steps 7 and 10, which applies the swap, $2$-opt, and shift operators with a first-improvement strategy. Our implementation is not particularly efficient, as it spends $O(n^{2})$ time on each operator.
4 Computational Results
We implemented Algorithms 13 in C++20 to evaluate their performance on the benchmark instances from Section 2.1. The code is freely available at https://github.com/fsoulignac/tsptw-m. We executed our experiments on a single thread of a laptop equipped with an AMD Ryzen 7 3700U CPU@2.32.3GHz and 66GB of physical RAM. To limit the impact of paging on the running time, we restricted the available RAM to 55GB.
4.1 Results for the TSPTW-M
For the TSPTW-M we set a time limit of $3$ minutes per instance. Although this time limit is quite restrictive for an exact method, it is sufficient for our purposes and helps prevent unnecessary computation time before encountering a memory limit exception. Furthermore, this time limit is representative of a typical use case, where an enumerative algorithm is applied for a short duration after computing a tight lower bound (see e.g. BaldacciMingozziRoberti2012; TilkIrnich2017; Lera-RomeroMirandaBrontSoulignac2022).
The results for the classical benchmarks and DaS are summarized in Table 1. We compare Algorithm 2 against the methods by Lera-RomeroMirandaBrontSoulignac2022 (Ler22), RudichCappartRousseau2023 (Rud23), and FontaineDibangoyeSolnon2023 (Fon23), which represent the state of the art for the TSPTW-M within exact methods. Ler22 and Fon23 were originally proposed for time-dependent versions of the TSPTW, and thus Lera-RomeroMirandaBrontSoulignac2022 and FontaineDibangoyeSolnon2023 did not report the results for the TSPTW-M. Similarly, RudichCappartRousseau2023 did not test Rud23 on the DaS benchmark. These missing results reported in Table 1 are taken from Fontaine2023. We note that the results in Fontaine2023 differ from those reported by RudichCappartRousseau2023 because Fontaine2023 executed all the algorithms on the same machine and with the same configuration. Our computer has a slightly higher performance for single thread execution according to www.cpubenchmark.net, and, unlike Fontaine2023, we did not disable turbo mode (we kept the ondemand governor within Ubuntu). Additionally, the time limit considered by Fontaine2023 was one hour per instance. These discrepancies do not affect our main conclusions, and we thus disregard them.
In Table 1, column # counts the number of instances in each benchmark set, columns $s$ indicate the number of solved instances, and columns $t_{s}$ report the average time for the solved instances, rounded to the nearest second. For Algorithm 2, we also include the maximum running time among the solved instances in column $m_{s}$. Additionally, we divide the results for Asc into Asc-s and Asc-l to separately report the results for instances with fewer than $50$ customers and those with $50$ or more customers, respectively. Table 1 reports the results for Rud23 with the unseeded version outside parentheses and the seeded version inside parentheses. The difference between the two is that the latter is initialized with a known solution. In this regard, Algorithm 2 should be considered unseeded. Similarly, Table 1 reports the results for Fon23 for two versions of the heuristic guiding the informed search. The results for FEA are reported outside parentheses, and those for MSA are reported inside parentheses.
Ler22 Rud23 Fon23 Algorithm 2
# $s^{*}$ $t_{s}^{*}$ $s$ $t_{s}$ $s^{*}$ $t_{s}^{*}$ $s$ $t_{s}$ $m_{s}$
Asc 50 48 21 48(48) 173(109) 50(49) 18(2) 50 13 152
Asc-s 32 32 21 152
Asc-l 18 18 0 0
DaS 125 125 485 112* 615* 125(125) 13(13) 125 2 7
Dum 135 135 39 133(135) 154(111) 135(135) 0(1) 135 0 1
Gen 130 91 524 122(125) 274(142) 117(110) 111(88) 130 0 1
Lan 70 70 0 70(70) 1(1) 70(70) 0(0) 70 0 0
Ohl 25 0 14(22) 1888(924) 20(20) 50(9) 25 1 6
Pes 27 22 203 26(27) 175(84) 25(27) 152(66) 25 16 164
Pot 30 25 160 28(28) 185(66) 27(29) 14(83) 25 7 80
Total 592 516 553(567) 569(565) 585
Table 1: Results of Algorithm 2 on the classical benchmarks. Entries marked with * were reported by Fontaine2023.
A quick look at Table 1 is enough to conclude that Algorithm 2 outperforms the state-of-the-art methods on the larger instances, solving the four instances that remained unsolved. Although this is strictly true, knowing the inner workings of Algorithm 2, we would never prefer it as a generic exact solver for the TSPTW-M. A deeper look at Table 1 raises an important concern: Algorithm 2 is unable to solve seven instances with 45 or fewer customers (and our unreported experiments show that memory is exhausted when the time limit is high enough). Furthermore, Algorithm 2 performs better on the larger instances in Asc than on the smaller ones. As strange as it may seem, RifkiSolnon2025 already noted that the ratio between the size of the time windows and the horizon could be more important than the number of customers when evaluating the efficiency of dynamic programming solvers for the TSPTW-M. Finally, we observe that Algorithm 2 is particularly fast on Lan, Dum, Gen, Ohl, and DaS, all of which follow the same creation procedure commonly applied in many machine learning methods to create “hard” training datasets (Section 2.1).
What we believe is happening is that Algorithm 2 takes advantage of some inherent bias in the benchmark instances resulting from the tight time windows, which yield a small search space of feasible solutions that have a well defined structure. Fon23 is another informed search that seems to have an advantage because of the structure of the instances, as depicted by the Ohl benchmark: the $20$ instances it solves are completed in just a few seconds, while Ler22 is unable to solve any of them within an hour! The difference is that Ler22 was designed to solve as many as possible of the small but challenging instances. As a result, it spends a large amount of time computing penalties on a relaxed problem before running its enumeration algorithm, and it uses a pure best-first search heuristic to guide its informed search, avoiding the extension of dominated labels. Finally, Rud23 is a generic algorithm based on multivalued decision diagrams that iteratively traverses enumeration trees of different widths to obtain lower and upper bounds to “peel” the tree. Being able to find feasible solutions quickly improves the peeling process.
To present the contrasting results, we report the performance of the algorithms for the Rif benchmark in Table 2. (The results for Ler22, Rud23, and Fon23 were taken from Fontaine2023). Again, Algorithm 2 is among the fastest for tight time windows ($\beta\in\{50,100\}$), but it is clearly the worst for the looser time windows ($\beta\in\{0,25\}$), even when compared against the generic method Rud23. (We remark that increasing the time limit does not help in these cases, as the memory limit is always reached.) The clear winner in robustness is Ler22, which was specifically designed for harder instances in terms of time windows.
$\beta$ $n$ Ler22 Rud23 Fon23 Algorithm 2
$s$ $t_{s}$ $s$ $t_{s}$ $s$ $t_{s}$ $s$ $t_{s}$ $m_{s}$
$100$ 2020 60 0 60 0 60 0 60 0 0
$30$ 60 0 60 1 60 0 60 0 0
$40$ 60 0 60 1 60 0 60 0 0
$50$ 2020 60 5 60 9 60 0 60 0 2
$30$ 60 129 58 105 60 1 59 5 125
$40$ 60 888 14 350 60 61 58 18 117
$25$ 2020 60 16 60 17 60 0 60 10 60
$30$ 60 643 27 361 60 36 0
$40$ 58 2318 0 47 1221 0
0 2020 60 59 60 25 60 0 60 35 177
$30$ 60 1376 10 434 60 177 0
$40$ 34 2837 0 5 2011 0
Total 692 469 652 506
Table 2: Results of Algorithm 2 on the Rif benchmark.
In our opinion, the main conclusion from the experiments in this section is that the classical benchmarks are no longer adequate for evaluating the TSPTW-M, as they can be “exploited” by a simple algorithm. Consequently, any method that incorporates a variant of Algorithm 2 in its preprocessing step would likely produce results that appear outstanding to an external reviewer.
4.2 Results for the TSPTW-D
For the TSPTW-D, we set a time limit of $30$ minutes per instance. Table 3 summarizes the results for the benchmarks Asc, DaS, Gen, Ohl, and Ohl-D. It includes the results of Algorithm 3 applied to the transport network $D$ (columns labeled Algorithm 3) and its reverse network $D^{-1}$ (columns labeled Algorithm 3-1), as the method is not symmetric, and the running times differ considerably between the two. For comparison, Table 3 also reports results from TilkIrnich2017 and Lera-RomeroMirandaBrontSoulignac2022.
We do not consider the benchmarks Pot and Pes because their travel times are rounded to $10^{-4}$, and Algorithm 3 cannot handle the large number of TSPTW-M instances in the range $10^{4}T]$. This also applies to the four unsolved instances in Asc (rbg021.xx for $x\in[\![6,9]\!]$), which were derived from rbg021 by increasing both the horizon and the deadline times in the time windows. Additionally, we exclude the benchmark Lan because its instances are trivial and were not tested by TilkIrnich2017 or Lera-RomeroMirandaBrontSoulignac2022. We also note that TilkIrnich2017; Lera-RomeroMirandaBrontSoulignac2022 did not evaluate their algorithms on the Gen instances with $n\geq 80$ and $\omega\in\{80,100\}$. It is worth mentioning that Lera-RomeroMirandaBrontSoulignac2022 solved all the instances in Pot, while TilkIrnich2017 solved all but three. However, both methods were tested with a time limit of 3 hours per instance.
Til22 Ler22 Algorithm 3 Algorithm 3-1
# $s$ $t_{s}$ $s$ $t_{s}$ $s$ $t_{s}$ $m_{s}$ $s$ $t_{s}$ $m_{s}$
Asc 50 50 55 47 8 46 9 295 45 3 60
DaS 125 125 2 7 125 2 8
Dum 135 135 0 1 135 0 1
Gen 130 97/115 633 114/115 667 123 30 1279 114 3 173
Ohl 25 25 15 216 25 3 27
Ohl-D 25 1 731 12 5159 20 10 59 25 92 1093
Table 3: Results of Algorithm 3 on the classical benchmarks.
Table 3 reinforces the conclusions from Section 4: Algorithm 3 (on either $D$ or $D^{-1}$) solves all but one of the classical instances with $50$ or more customers, including the 14 instances that were not solved by TilkIrnich2017 or Lera-RomeroMirandaBrontSoulignac2022. Furthermore, Algorithm 3 outperforms both methods on the larger instances with higher values of $n$.
By design, one might expect Algorithm 3 to be much less efficient than Algorithm 2, as it solves $\Theta(T)$ instances of the TSPTW-M in the worst case. However, as implied by Table 3, only a small number of these instances are actually passed to Algorithm 2 for solving in the benchmark instances. Most feasible departure times $p$ in Loop 611 are either solved by the local search heuristic or ignored because of the best-known solution $R^{*}$. Moreover, $\operatorname{last}_{p}$ is typically small in the largest instances, which helps reduce the number of instances solved by Algorithm 2.
The main conclusion is that the TSPTW-D should be evaluated on “harder” instances, especially those with wider time windows, to better assess the performance of algorithms in more challenging scenarios.
5 Conclusions
We presented a simple informed search method for the TSPTW-M, which excels in solving all classical benchmark instances with 50 or more customers in under 10 seconds each. However, it struggles to solve even the smallest instances when their time windows are sufficiently loose. By running the algorithm as a preprocessing step for a limited amount of time, any method can solve these instances, meaning they should no longer be used for benchmarking purposes.
An important question that arises from this work is how valid and accurate are the conclusions drawn from methods that have been tested exclusively on these classical benchmark instances? Are these methods truly efficient in a general setting, or are they, perhaps unconsciously, exploiting some inherent bias in these benchmarks? Since these instances can be "hacked" by simpler methods, it is crucial to examine how generalizable the reported results are. Exploring this issue further could help ensure that conclusions drawn from these benchmarks are not only valid for the specific instances tested but also robust across a wider range of real-world problem scenarios.
Our method performs particularly well on the benchmarks Lan, Dum, Gen, Ohl, and DaS, all of which share the common strategy of defining a time window $x-\omega,x+\omega]$ for a vertex visited at time $x$ in a given route. Here, $\omega$ is relatively small when compared to $n$ and $T$. Due to the simplicity of this approach, it is commonly used to generate training datasets for many machine learning methods, which can lead to overfitting and biased conclusions. As noted by RifkiSolnon2025, a simple solution to make these instances more challenging for an informed search algorithm like ours is to extend the time windows. The varying levels of difficulty among these generated instances provide a better foundation for drawing deeper conclusions.
Given the exceptional performance of our method on the larger benchmark instances, it is not surprising that a trivial procedure of repeatedly running the solver across the entire horizon is sufficient to solve the TSPTW-D on most of the instances. The issue is that many of these instances are easy even for the TSPTW-D. Therefore, these instances should also be avoided when evaluating methods for the TSPTW-D. What remains to be explored is whether a similar informed search can be applied to solve the TSPTW-TT on these instances. Nonetheless, we believe it would be wise to evaluate and train future methods on instances with varying levels of time window tightness, even for the TSPTW-TT.